Ask
your own question, for FREE!
Mathematics
12 Online
There is a 70% chance of heavy fog and a 30% chance of light drizzle. A softball game has an 80% chance of being cancelled for a light drizzle and a 30% chance of being cancelled for a heavy for when there is no drizzle. Find the probability the game will be cancelled.
Still Need Help?
Join the QuestionCove community and study together with friends!
Prob of heavy fog , H = 0.7 Prob of light drizzle, L = 0.3 Prob of game cancelled during H, X= 0.3 Prob of game cancelled during L, Y = 0.8 Prob of game cancelled = XH + YL = 0.21 + 0.24 = 0.45
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
7 hours ago
0 Replies
0 Medals
3 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
1 day ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal